### v2

# Some remarks on Gorenstein projective precovers

Abstract.
In this paper, we prove that for a $n$-perfect ring $R$ various classes of relative Gorenstein projective $R$-modules are special precovering, among them including the Gorenstein projectives and Ding projectives. We got a family of hereditary and complete cotorsion pairs and hereditary Hovey triples that comes from relative Gorenstein projective $R$-modules. We also introduce a new class of Gorenstein projective $R$-modules relative to the Auslander class $\mathcal{A}_{C}(R)$ of a semidualizing $(R,S)$-bimodule ${}_{R}C_{S}$.
2010 Mathematics Subject Classification. Primary 18G10, 18G20, 18G25. Secondary 16E10.
Key Words: Gorenstein projective, Contravariantly finite, Hovey triple, Auslander class, Duality pair
1. Introduction
A central question in homological theory of modules is to understand the size of the homological dimensions which are associated with rings and modules. In particular the famous second finitistic conjecture is established from the question of how large can the projective dimension of modules over a ring be?. To address this concern consider a ring $R$ and $\mathcal{P}(R)^{<\infty}_{fin}$ the class of $R$-modules finitely generated of finite projective dimension, the Second Finitistic Dimension Conjeture establish that $fpd(R):=\sup\{\mathrm{pd}_{R}(M)|M\in\mathcal{P}(R)^{<\infty}_{fin}\}$ is finte.
It is well-known that the understanding of the class $\mathcal{P}(R)^{<\infty}_{fin}$ and $\mathcal{P}(R)^{<\infty}$ provides insight to finitistic dimension conjectures as was proved by Auslander and Reiten in [1] who proved that contravariant finiteness of the class [^1] In this paper we use precovering as synonym of contravariantly finite. $\mathcal{P}(R)^{<\infty}_{fin}$, referred to as the Auslander–Reiten condition from now on, is a sufficient condition for validity of the second finitistic dimension conjecture over an Artin algebra. More recently it is proved by P. Moradifar and J. Šaroch [21] that contravariant finiteness of the class $\mathcal{GP}^{<\infty}_{fin}$ (finitely generated modules of finite Gorenstein projective dimension) implies validity of the second finitistic dimension conjecture over left artinian rings. Furthermore is proved that that contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$ implies contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, over rings where $\mathcal{GP}(R)_{fin}$ (the class of finitely generated Gorenstein projective modules) is contravariantly finite and the converse holds for Artin algebras. In a more general context S. Estrada, A. Iacob and K. Yeomans have proven that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$ provided that the ring $R$ be GF-closed, right coherent and left $n$-perfect [13, Theorem 2]. And more recently, a relationship has been found between the existence of Gorenstein projective precovers and finitely presented $R$-modules with the Second Finitistic Dimension Conjeture [11, Theorem 5]. Thus, in view of the importance contravariant finiteness of the class $\mathcal{GP}(R)$, we study when such class and their generalizations are special precovering on $\mathrm{Mod}(R)$. In this paper we prove in Proposition 3.2 and Corollary 4.8 that when $R$ is left $n$-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$ the classes of Gorenstein projectives $\mathcal{GP}(R)$ and Ding projectives $\mathcal{DP}(R)$ are special precovering. Also in Proposition 3.4 we prove that the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ is special precovering provided that $R$ is left $n$-perfect. We then turn our attention to the classes of Gorenstein projective $R$-modules that come from a duality pair $(\mathcal{L},\mathcal{A})$ and we prove in Proposition 4.9 that for a complete duality pair over a $n$-perfect ring $R$, the class of Gorenstein projectives $\mathcal{GP}_{\mathcal{L}}(R)$ relative to this pair $(\mathcal{L},\mathcal{A})$ is special precovering. As an application, we recover in Proposition 4.11 the main result of [13]. In Theorem 4.14 we get families of cotorsion pairs and in Corollary 4.15 families of Hovey triples. At the end in Proposition 4.16 we declare an application of the developed theory on a new class of Gorenstein projective $R$-modules.
2. Preliminaries
In what follows, we shall work with categories of modules over an associative ring $R$ with identity. By $\mathrm{Mod}(R)$ and $\mathrm{Mod}(R^{\rm op})$ we denote the categories of left and right $R$-modules.
Projective, injective and flat $R$-modules will be important to present some definitions, remarks and examples. The classes of projective left and right $R$-modules will be denoted by $\mathcal{P}(R)$ and $\mathcal{P}(R^{\rm op})$, respectively. Similarly, we shall use the notations $\mathcal{I}(R)$, $\mathcal{I}(R^{\rm op})$, $\mathcal{F}(R)$ and $\mathcal{F}(R^{\rm op})$ for the classes of injective and flat modules in $\mathrm{Mod}(R)$ and $\mathrm{Mod}(R^{\rm op})$, respectively.
Concerning functors defined on modules, $\mathrm{Ext}^{i}_{R}(-,-)$ denotes the right $i$-th derived functor of $\mathrm{Hom}_{R}(-,-)$. If $M\in\mathrm{Mod}(R^{\rm op})$ and $N\in\mathrm{Mod}(R)$, $M\otimes_{R}N$ denotes the tensor product of $M$ and $N$. Recall the construction of this tensor products defines a bifunctor $-\otimes_{R}-\colon\mathrm{Mod}(R^{\rm op})\times\mathrm{Mod}(R)\longrightarrow\mathrm{Mod}(\mathbb{Z})$, where $\mathrm{Mod}(\mathbb{Z})$ is the category of abelian groups.
Orthogonality
Let $\mathcal{X}\subseteq\mathrm{Mod}(R)$, $i\geq 1$ be a positive integer and $N\in\mathrm{Mod}(R)$. The expression $\mathrm{Ext}^{i}_{R}(\mathcal{X},N)=0$ means that $\mathrm{Ext}^{i}_{R}(X,N)=0$ for every $X\in\mathcal{X}$. Moreover, $\mathrm{Ext}^{i}_{R}(\mathcal{X,Y})=0$ if $\mathrm{Ext}^{i}_{R}(\mathcal{X},Y)=0$ for every $Y\in\mathcal{Y}$. The expression $\mathrm{Ext}^{i}_{R}(N,\mathcal{Y})=0$ has a similar meaning. Moreover, by $\mathrm{Ext}^{\geq 1}_{R}(M,N)=0$ we mean that $\mathrm{Ext}^{i}_{R}(M,N)=0$ for every $i\geq 1$. One also has similar meanings for $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X},N)=0$, $\mathrm{Ext}^{\geq 1}_{R}(N,\mathcal{Y})=0$ and $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$. We can also replace $\mathrm{Ext}$ by $\mathrm{Tor}$ in order to obtain the notations for $\mathrm{Tor}$-orthogonality. The right orthogonal complements of $\mathcal{X}$ will be denoted by
$\mathcal{X}^{\perp_{i}}$ :=$:=\{M\in\mathrm{Mod}(R){\rm\ :\ }\mathrm{Ext}^{i}_{R}(\mathcal{X},M)=0\}$ and $\mathcal{X}^{\perp}$ $:=\bigcap_{i\geq 1}\mathcal{X}^{\perp_{i}}.$
The left orthogonal complements, on the other hand, are defined similarly.
Relative homological dimensions
There are homological dimensions defined in terms of extension functors. Let $M\in\mathrm{Mod}(R)$ and $\mathcal{X},\mathcal{Y}\subseteq\mathrm{Mod}(R)$. The injective dimensions of $M$ and $\mathcal{Y}$ relative to $\mathcal{X}$ are defined by
$\mathrm{id}_{\mathcal{X}}(M)$ $:=\inf\{m\in\mathbb{Z}_{\geq 0}\text{ : }\mathrm{Ext}_{R}^{\geq m+1}(\mathcal{X},M)=0\}$ and $\mathrm{id}_{\mathcal{X}}(\mathcal{Y})$ $:=\sup\{\mathrm{id}_{\mathcal{X}}(Y)\text{ : }Y\in\mathcal{Y}\}.$
In the case where $\mathcal{X}=\mathrm{Mod}(R)$, we write
$\mathrm{id}_{\mathrm{Mod}(R)}(M)$ $=\mathrm{id}(M)$ and $\mathrm{id}_{\mathsf{Mod}(R)}(\mathcal{Y})$ $=\mathrm{id}(\mathcal{Y})$
for the (absolute) injective dimensions of $M$ and $\mathcal{Y}$. Dually we can define the relative dimensions $\mathrm{pd}_{\mathcal{X}}(M)$, $\mathrm{pd}_{\mathcal{X}}(\mathcal{Y})$ and $\mathrm{pd}(M)$, $\mathrm{pd}(\mathcal{Y})$.
By an $\mathcal{X}$-resolution of $M$ we mean an exact complex
$\cdots\to X_{m}\to X_{m-1}\to\cdots\to X_{1}\to X_{0}\to M\to 0$
with $X_{k}\in\mathcal{X}$ for every $k\in\mathbb{Z}_{\geq 0}$. If $X_{k}=0$ for $k>m$, we say that the previous resolution has length $m$. The resolution dimension relative to $\mathcal{X}$ (or the $\mathcal{X}$-resolution dimension) of $M$ is defined as the value
$\mathrm{resdim}_{\mathcal{X}}(M):=\min\{m\in\mathbb{Z}_{\geq 0}\ \mbox{ : }\ \text{there exists an $\mathcal{X}$-resolution of $M$ of length $m$}\}.$
Moreover, if $\mathcal{Y}\subseteq\mathrm{Mod}(R)$ then
$\mathrm{resdim}_{\mathcal{X}}(\mathcal{Y}):=\sup\{\mathrm{resdim}_{\mathcal{X}}(Y)\ \mbox{ : }\ \text{$Y\in\mathcal{Y}$}\}$
defines the resolution dimension of $\mathcal{Y}$ relative to $\mathcal{X}$. The classes of objects with bounded (by some $n\geq 0$) and finite $\mathcal{X}$-resolution dimensions will be denoted by
$\mathcal{X}^{\wedge}_{n}$ :=$:=\{M\in\mathrm{Mod}(R)\text{ : }\mathrm{resdim}_{\mathcal{X}}(M)\leq n\}$ and $\mathcal{X}^{\wedge}$ $:=\bigcup_{n\geq 0}\mathcal{X}^{\wedge}_{n}.$
Dually, we can define $\mathcal{X}$-coresolutions and the coresolution dimension of $M$ and $\mathcal{Y}$ relative to $\mathcal{X}$ (denoted $\mathrm{coresdim}_{\mathcal{X}}(M)$ and $\mathrm{coresdim}_{\mathcal{X}}(\mathcal{Y})$). We also have the dual notations $\mathcal{X}^{\vee}_{n}$ and $\mathcal{X}^{\vee}$ for the classes of $R$-modules with bounded and finite $\mathcal{X}$-coresolution dimension.
Approximations
Given a class $\mathcal{X}$ of left $R$-modules and $M\in\mathrm{Mod}(R)$, recall that a morphism $\varphi\colon X\to M$ with $X\in\mathcal{X}$ is an $\mathcal{X}$-precover of $M$ if for every morphism $\varphi^{\prime}\colon X^{\prime}\to M$ with $X^{\prime}\in\mathcal{X}$, there exists a morphism $h\colon X^{\prime}\to X$ such that $\varphi^{\prime}=\varphi\circ h$. An $\mathcal{X}$-precover $\varphi$ is special if $\mathrm{CoKer}(\varphi)=0$ and $\mathrm{Ker}(\varphi)\in\mathcal{X}^{\bot_{1}}$.
A class $\mathcal{X}\subseteq\mathrm{Mod}(R)$ is (pre)covering if every left $R$-module has an $\mathcal{X}$-(pre)cover. Dually, one has the notions of (pre)envelopes and (pre)enveloping and special (pre)enveloping classes.
Cotorsion pairs
Two classes $\mathcal{X,Y}\subseteq\mathrm{Mod}(R)$ of left $R$-modules for a cotorsion pair $(\mathcal{X,Y})$ if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$.
A cotorsion pair $(\mathcal{X,Y})$ is:
  • Complete if $\mathcal{X}$ is special precovering, that is, for every $M\in\mathrm{Mod}(R)$ there is a short exact sequence $0\to Y\to X\to M\to 0$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$; or equivalently, if $\mathcal{Y}$ is special preenveloping.
  • Hereditary if $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$; or equivalently, if $\mathcal{X}$ is resolving (meaning that $\mathcal{X}$ is closed under extensions and kernels of epimorphisms, and contains the projective left $R$-modules) or $\mathcal{Y}$ is coresolving.
Note that if $(\mathcal{X,Y})$ is a hereditary cotorsion pair, then $\mathcal{X}={}^{\perp}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp}$.
Duality pairs
The notion of duality pair was introduced by Holm and Jørgensen in [18]. Two classes $\mathcal{L}\subseteq\mathrm{Mod}(R)$ and $\mathcal{A}\subseteq\mathrm{Mod}(R^{\rm op})$ of left and right $R$-modules form a duality pair $(\mathcal{L,A})$ in $\mathrm{Mod}(R)$ if:
  1. (1)
    $L\in\mathcal{L}$ if, and only if, $L^{+}:=\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Q}/\mathbb{Z})\in\mathcal{A}$.
  2. (2)
    $\mathcal{A}$ is closed under direct summands and finite direct sums.
One has a similar notion of duality pair in the case where $\mathcal{L}$ is a class of right $R$-modules, and $\mathcal{A}$ is a class of left $R$-modules.
A duality pair $(\mathcal{L,A})$ is called:
  • (co)product-closed if $\mathcal{L}$ is closed under (co)products.
  • perfect if it is coproduct closed, $\mathcal{L}$ is closed under extensions and contains $R$ (regarded as a left $R$-module).
  • complete if $(\mathcal{A},\mathcal{L})$ is also a duality pair and $(\mathcal{L},\mathcal{A})$ has all the properties required to be a perfect duality pair.
3. Gorenstein projective precovers
We recall that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ consist of cycles of exact complexes of left projective $R$-modules which remains exact after applying the functor $\mathrm{Hom}_{R}(-,P)$ for all $P\in\mathcal{P}(R)$. Also the class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$ consist of cycles of exact complexes of left flat $R$-modules which remains exact after applying the funtor $I\otimes_{R}-$, for all $I\in\mathcal{I}(R^{\mathrm{op}})$.
From Šaroch and Št’ovíchěk’s [23] we know now that every ring is GF-closed [^2] This means that $\mathcal{GF}(R)$ is closed by extensions for all ring $R$.. Thus, is enough that the ring $R$ will be right coherent and left $n$-perfect [^3] This means that for the ring $R$ and $n\geq 0$ the containment $\mathcal{P}(R)\subseteq\mathcal{F}(R)^{\wedge}_{n}$ is given. to obtain $\mathcal{GP}(R)$ special precovering. In what follows we will see that such result remains true under other conditions less known and that the above result can be obtained for classes of relative Gorenstein projective $R$-modules.
Proposition 3.1.
Consider $R$ a left $n$-perfect ring. Then, every Gorenstein flat $R$-module has a $\mathcal{GP}(R)$-resolution at most $n$. That is, the containment $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$ is given.
Proof.
Indeed, from [12, Proposition 3.1 (1)] the condition $\mathrm{Ext}_{R}^{i}(G,F)=0$ for all $G\in\mathcal{GF}(R)$ and $F\in\mathcal{F}\cap\mathcal{F}^{\bot}$ is true without conditions over the ring. Thus by using [13, Proposition 2] we conclude that $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$. ∎
We have here our first main result.
Proposition 3.2.
Let $R$ be a ring left $n$-perfect and such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. Then the class $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$.
Proof.
We will verify the conditions of [13, Theorem 1]. From [27, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$, since $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. While the containment $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$ it follows from Proposition 3.1. ∎
This situation also arises for the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ introduced by Šaroch and Št’ovíchěk [23]. Such class $\mathcal{P}\mathcal{GF}(R)$ consist of cycles of exact complexes of left projective $R$-modules which remains exact after applying the funtor $I\otimes_{R}-$, for all $I\in\mathcal{I}(R^{\mathrm{op}})$. To see how this happens, we need a useful tool, which is more extensive than the one presented in [9, Proposition 3.1]. The notation $P\mathrm{Gfd}(M)$ means $\mathrm{resdim}_{\mathcal{P}\mathcal{GF}}(M)$, called the projectively coresolved Gorenstein flat dimension of $M$.
Proposition 3.3.
Let $R$ be a ring and consider $\mathcal{P}\mathcal{GF}(R)$ the class of projectively coresolved Gorenstein flat $R$-modules. Then for all $M\in\mathrm{Mod}(R)$ with $P\mathrm{Gfd}(M)=n<\infty$, there are exact sequences in $\mathrm{Mod}(R)$
$0\to K\to X\stackrel{{\scriptstyle\varphi}}{{\to}}M\to 0,$
with $\mathrm{pd}(K)=n-1$ and $X\in\mathcal{P}\mathcal{GF}(R)$, where $\varphi:X\to M$ is an $\mathcal{P}\mathcal{GF}(R)$-precover and $K\in\mathcal{P}\mathcal{GF}(R)^{\perp}$.
And
$0\to M\stackrel{{\scriptstyle\varphi^{\prime}}}{{\to}}H\to X^{\prime}\to 0,$
with $\mathrm{pd}(H)=n$ and $X^{\prime}\in\mathcal{P}\mathcal{GF}(R).$ Where $\varphi^{\prime}:M\to H$ is an $\mathcal{P}(R)^{\wedge}$-preenvelope and $X^{\prime}\in{}^{\perp}(\mathcal{P}(R)^{\wedge})$.
Proof.
Since $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GP}(R)$ [23, Theorem 4.4] and the pair $(\mathcal{GP}(R),\mathcal{P}(R))$ is a strong left Frobenius pair then the pair $(\mathcal{P}\mathcal{GF}(R),\mathcal{P}(R))$ forms also a strong left Frobenius pair, since $\mathcal{P}\mathcal{GF}(R)$ is projective resolving closed under direct sums and direct summands. Thus, the result it follows from apply [6, Theorem 2.8]. ∎
We have here our second main result.
Proposition 3.4.
Let $R$ be a left $n$-perfect ring, then the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ is special precovering on $\mathrm{Mod}(R)$.
Proof.
Actually we know that $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GF}(R)$, without conditions on $R$. Now, from [9, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}_{n}$. We will use this in the proof, as follows. Let $X\in\mathrm{Mod}(R)$, by [14, Theorem 3.4] we know that $(\mathcal{GF}(R),\mathcal{GF}(R)^{\bot})$ is a hereditary and complete cotorsion pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}(R)$ and $Y\in\mathcal{GF}(R)^{\bot}$. From the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GF}(R)$, we get $\mathcal{GF}(R)^{\bot}\subseteq\mathcal{P}\mathcal{GF}(R)^{\bot}$. Since $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GP}(R)^{\wedge}_{n}$. By Proposition 3.3 for $N\in\mathcal{GF}(R)$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{P}\mathcal{GF}(R)$ and $W\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. We can construct the following p.b digram
[…]
to obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. This implies that $A\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. Therefore we obtain the sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{P}\mathcal{GF}(R)$ and $A\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. Therefore $T\to X$ is a special $\mathcal{P}\mathcal{GF}(R)$-precover. ∎
4. Generalized Gorenstein flat modules
In what follows, we shall consider classes $\mathcal{X}\subseteq\mathrm{Mod}(R)$ and $\mathcal{Y}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. The following definition of Gorenstein flat $R$-modules relative to a pair $(\mathcal{X},\mathcal{Y})$ comes from [28, Definition 2.1].
Definition 4.1.
An $R$-module $M$ is Gorenstein $(\mathcal{X},\mathcal{Y})$-flat if there exists an exact and $(\mathcal{Y}\otimes_{R}-)$-acyclic complex $X_{\bullet}\in\mathrm{Ch}(\mathcal{X})$ such that $M\simeq Z_{0}(X_{\bullet})$. By $(\mathcal{Y}\otimes_{R}-)$-acyclic we mean that $Y\otimes_{R}X_{\bullet}$ is an exact complex of abelian groups for every $Y\in\mathcal{Y}$. The class of Gorenstein $(\mathcal{X,Y})$-flat $R$-modules will be denoted by $\mathcal{GF}_{(\mathcal{X,Y})}$.
There are several types of classes Gorenstein $(\mathcal{X},\mathcal{Y})$-flat that have been presented in recent literature. In [23, §4] is presented by J. Šaroch and J. Št’ovíček, the class of projectively coresolved Gorenstein flat $R$-modules, here denoted $\mathcal{P}\mathcal{GF}(R)$, later generalized and extensively studied by S. Estrada, A. Iacob and M. A. Pérez in [12, Definition 2.6]. This class is called projectively coresolved Gorenstein $\mathcal{B}$-flat and here denoted $\mathcal{P}\mathcal{GF}_{\mathcal{B}}$. In such paper also is studied the class $\mathcal{GF}_{\mathcal{B}}$ (in the notation of Definition 4.1 is $\mathcal{GF}_{(\mathcal{F},\mathcal{B})}$ where the subscript $\mathcal{F}$ denotes $\mathcal{F}(R)$) with $\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ is sometimes a semi-definable class [12, Definition 2.7].
In a different environment is defined by J. Gillespie in [16] for a complete duality pair $(\mathcal{L},\mathcal{A})$ the class of Gorenstein $(\mathcal{L},\mathcal{A})$-flat $R$-modules, here denoted by $\mathcal{GF}_{\mathcal{A}}$, and is studied [16, §5] some of their model structures. In such paper also is defined the class of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules here denoted by $\mathcal{GP}_{\mathcal{L}}$ and the class of Gorenstein $(\mathcal{L},\mathcal{A})$-injective $R$-modules here denoted by $\mathcal{GI}_{\mathcal{A}}$. We see that when $(\mathcal{L},\mathcal{A})$ is a duality pair the classes $\mathcal{GP}_{\mathcal{L}}$, $\mathcal{GI}_{\mathcal{A},}$, $\mathcal{GF}_{\mathcal{A}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ have an interesting interaction between them (see [4], [12], [16], [17], [26]). In fact the classes $\mathcal{GP}_{\mathcal{L}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ agree sometimes. This is true for a symmetric duality pair $(\mathcal{L},\mathcal{A})$, from [7, Theorem A.6.] [^4] Note that for D. Bravo et. al, the notion of duality pair differs from the one presented here. since the class $\mathcal{L}$ is closed by pure quotients [18, Theorem 3.1]. This fact had been mentioned previously by J. Gillespie and A. Iacob in [17, Corollary 14]. We declare this result as follows.
Proposition 4.2.
Given a symmetric duality pair $(\mathcal{L},\mathcal{A})$ in $\mathrm{Mod}(R)$ the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ of projectively coresolved Gorenstein $\mathcal{A}$-flat $R$-modules coincide with the class $\mathcal{GP}_{\mathcal{L}}$ of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules.
When for a duality pair $(\mathcal{L},\mathcal{A})$ the class $\mathcal{A}$ is semi-definable then $\mathcal{GP}_{\mathcal{L}}$ is special precovering in the whole category $\mathrm{Mod}(R)$, since [12, Theorem 2.13] the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ is the left part of a complete and hereditary cotorsion pair. A duality pair with such property is the definable pair mentioned in [17, Example 12] and denoted $(<\mathcal{F}(R)>,<\mathcal{I}(R^{\mathrm{op}})>)$. That is, the class of Gorenstein $(\mathcal{P}(R),<\mathcal{F}(R)>)$-projective $R$-modules is special precovering in $\mathrm{Mod}(R)$. We state this as follows.
Proposition 4.3.
Consider the definable duality pair $(<\mathcal{F}(R)>,<\mathcal{I}(R^{\mathrm{op}})>)$ in $\mathrm{Mod}(R)$, then the class $\mathcal{GP}_{<\mathcal{F}(R)>}$ is the left part of a complete and hereditary cotorsion pair.
For each $n\geq 2$ there is a duality pair $(\mathcal{FP}_{n}\mbox{-}Flat(R),\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}}))$, which consists of the $FP_{n}$-flat left R-modules and $FP_{n}$-injective right $R$-modules [8, Theorems 5.5 & 5.6]. The right part $\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}})$ is definable [12, Example 2.21. (3)]. The classes of Gorenstein $R$-modules that comes from these duality pairs have been studied in [20] by A. Iacob, proving in a manner similar to that used in Proposition 4.3 that $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ is the left part of a complete and hereditary cotorsion pair, and thus is a class special precovering. From here, we have the following result.
Proposition 4.4.
If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given then the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering.
Proof.
Note that we have $\mathcal{P}(R)\subseteq\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [4, Proposition 6.7] the classes $\mathcal{GP}(R)$ and $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ are the same. Now from [20, Theorem 3.8] we get that $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair. ∎
The purpose of this work is to continue with the study of the conditions over a duality pair $(\mathcal{L},\mathcal{A})$ that implies that $\mathcal{GP}_{\mathcal{L}}$ will be special precovering, this conditions will be different to the notion of semidefinable for the class $\mathcal{A}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. For do this we study interactions between the classes of relative Gorentein flat and relative Gorenstein projective associated to a duality pair $(\mathcal{L},\mathcal{A})$. We begin by recalling some facts and a definition adapted to our setting.
Definition 4.5 (GP-admissible pair).
[5, Definition 3.1]. A pair $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}(R)\times\mathrm{Mod}(R)$ is GP-admissible if satisfies the following conditions:
  1. (1)
    $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$.
  2. (2)
    For every $A\in\mathrm{Mod}(R)$ there is an epimorphism $X\to A$ with $X\in\mathcal{X}$.
  3. (3)
    $\mathcal{X}$ and $\mathcal{Y}$ are closed under finite coproducts.
  4. (4)
    $\mathcal{X}$ is closed under extensions.
  5. (5)
    $\mathcal{X}\cap\mathcal{Y}$ is a relative cogenerator in $\mathcal{X}$, that is, for every $X\in\mathcal{X}$ there is an exact sequence $0\to X\to W\to X^{\prime}\to 0$ with $X^{\prime}\in\mathcal{X}$ and $W\in\mathcal{X}\cap\mathcal{Y}$.
Remark 4.6.
  • (i)
    We see from [12, Corollary 2.20] that for $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ and when $\mathcal{GF}_{\mathcal{B}}$ is closed under extensions, then $(\mathcal{GF}_{\mathcal{B}},\mathcal{GF}_{\mathcal{B}}^{\bot_{1}})$ is a hereditary and complete cotorsion pair in $\mathrm{Mod}(R)$. Furthermore from [12, Proposition 3.1] we have that $\mathcal{GF}_{\mathcal{B}}\cap\mathcal{GF}_{\mathcal{B}}^{\bot}=\mathcal{F}(R)\cap\mathcal{F}(R)^{\bot}$.
  • (ii)
    Note that when $(\mathcal{L},\mathcal{A})$ is a perfect duality pair, then $\mathcal{P}(R)\subseteq\mathcal{L}$, $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{A}$ and the pair $(\mathcal{P}(R),\mathcal{L})$ is GP-admissible in $\mathrm{Mod}(R)$ (see [4, §3]).
  • (iii)
    For a complete duality pair $(\mathcal{L},\mathcal{A})$ we have from [17, Corollary 43] that $(\mathcal{GF}_{\mathcal{A}},\mathcal{GF}_{\mathcal{A}}^{\bot_{1}})$ is a perfect cotorsion pair. In consequence the class $\mathcal{GF}_{\mathcal{A}}$ is closed by extensions and from $\mathrm{(i)}$ and $\mathrm{(ii)}$ also is an hereditary cotorsion pair.
We are ready to prove the third main result, which is a generalization of [13, Main Result].
Theorem 4.7.
Let $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}\times\mathrm{Mod}(R^{\mathrm{op}})$ with the following properties.
  • (i)
    $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{Y}$.
  • (ii)
    The pair $(\mathcal{P}(R),\mathcal{X})$ is GP-admissible.
  • (iii)
    The class $\mathcal{GF}_{\mathcal{Y}}$ is closed under extensions.
Assume that the inclusions $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ are given. Then the class $\mathcal{GP}_{\mathcal{X}}$ is special precovering in $\mathrm{Mod}(R)$.
Proof.
Take $X\in\mathrm{Mod}(R)$. From Remark 4.6 (i), we have that $(\mathcal{GF}_{\mathcal{Y}},\mathcal{GF}_{\mathcal{Y}}^{\bot})$ is a hereditary and complete cotorsión pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}_{\mathcal{Y}}$ and $Y\in\mathcal{GF}_{\mathcal{Y}}^{\bot}$. From the containment $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}$, we get $\mathcal{GF}_{\mathcal{Y}}^{\bot}\subseteq\mathcal{GP}_{\mathcal{X}}^{\bot}$. Since $\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ and $(\mathcal{P}(R),\mathcal{X})$ is a GP-admissible pair, then from [5, Theorem 4.1 (a)] for $N\in\mathcal{GF}_{\mathcal{Y}}$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{GP}_{\mathcal{X}}$ and $W\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. We can construct the following p.b digram
[…]
and obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{GP}_{\mathcal{X}}^{\bot}$, this implies that $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. Therefore we obtain the sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{GP}_{\mathcal{X}}$ and $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. This is, $T\to X$ is a special $\mathcal{GP}_{\mathcal{X}}$-precover. ∎
Corollary 4.8.
Let $R$ be a left n-perfect ring and such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. Then, the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering in $\mathrm{Mod}(R)$.
Proof.
We will verify the conditions of the Theorem 4.7 with the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$. To this end note that $(\mathcal{P}(R),\mathcal{F}(R))$ is a GP-admissible pair and $\mathcal{GP}_{\mathcal{F}}$ is precisely the class of Ding-projective $R$-modules $\mathcal{DP}(R)$, while $\mathcal{GF}_{\mathcal{I}(R^{\mathrm{op}})}$ is the usual class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$. From [27, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ and always is true that $\mathcal{DP}(R)\subseteq\mathcal{GP}(R)$, therefore $\mathcal{DP}(R)\subseteq\mathcal{GF}(R)$. Also, from [9, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}_{n}$ and by [17, Corollary 14] the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{DP}(R)$ is always true. All this give us $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}\subseteq\mathcal{DP}(R)^{\wedge}$. ∎
Until now, from Proposition 4.2 we know that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality the containment $\mathcal{GP}_{\mathcal{L}}\subseteq\mathcal{GF}_{\mathcal{A}}$ is given. By Remark 4.6 if $(\mathcal{L},\mathcal{A})$ is perfect, then are fulfilled the conditions $\mathrm{(i),(ii)}$ in the previous Theorem 4.7. While if $(\mathcal{L},\mathcal{A})$ is complete then $\mathrm{(iii)}$ in Theorem 4.7 is also true. In the present generality, we have the following result.
Proposition 4.9.
Let $(\mathcal{L},\mathcal{A})$ be a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect ring. Then the class $\mathcal{GP}_{\mathcal{L}}$ is special precovering on $\mathrm{Mod}(R)$.
Proof.
Since $(\mathcal{L},\mathcal{A})$ complete, also is symmetric and perfect. Thus, the conditions in Theorem 4.7 are fulfilled, except for the containment $\mathcal{GF}_{\mathcal{A}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{L}}$. To prove such a containment we must assume that $R$ is $n$-perfect, that is $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$.
Take $G\in\mathcal{GF}_{\mathcal{A}}$, then there is an exact complex $\mathbf{N}$ of flat $R$-modules $(\mathcal{A}\otimes_{R}-)$-acyclic with $G=Z_{0}(\mathbf{N})$. Consider a partial projective resolution of $\mathbf{N}$
$0\to\mathbf{C}\to\mathbf{P}_{n-1}\xrightarrow{d_{n-1}}\mathbf{P}_{n-2}\xrightarrow{d_{n-2}}\cdots\to\mathbf{P}_{1}\xrightarrow{d_{1}}\mathbf{P}_{0}\xrightarrow{d_{0}}\mathbf{N}\to 0,$
where $\mathbf{C}$ is not projective (but will be an exact complex). From this, for each $j$ we have the exact sequence $0\to C_{j}\to P_{n-1,j}\to\cdots\to P_{0,j}\to N_{j}\to 0$, with every $P_{i,j}\in\mathcal{P}(R)$, and since $N_{j}\in\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$ it follows that $C_{j}$ is projective for all $j$. Also we have the exact sequence $0\to\mathrm{Ker}(d_{0})\to\mathbf{P}_{0}\to\mathbf{N}\to 0$, with $\mathrm{Ker}(d_{0})$ exact, since $\mathbf{P}_{0}$ and $\mathbf{N}$ are exact. For each $A\in\mathcal{A}$ by assumption $A\otimes_{R}\mathbf{N}$ is acyclic, and since $\mathbf{P}_{0}$ is an projective complex then $A\otimes_{R}\mathbf{P}_{0}$ is acyclic, this implies that $A\otimes_{R}\mathrm{Ker}(d_{0})$ is acyclic. Repeating this procedure we obtain that $\mathbf{C}$ is an exact complex and $A\otimes_{R}\mathbf{C}$ is acyclic for all $A\in\mathcal{A}$. Therefore $\mathbf{C}$ is an exact complex of projectives $(\mathcal{A}\otimes_{R}-)$-acyclic, i.e. $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}$. Note that for each $j$ we have the exact sequence
$0\to Z_{j}(\mathbf{C})\to Z_{j}(\mathbf{P}_{n-1})\to\cdots Z_{j}(\mathbf{P}_{0})\to Z_{j}(\mathbf{N})\to 0$
with $Z_{j}(\mathbf{P}_{i})\in\mathcal{P}(R)$ for each $j$ and where $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}=\mathcal{GP}_{\mathcal{L}}$ by Proposition 4.2. Thus for $j=0$ we get the desired $\mathcal{GP}_{\mathcal{L}}$-resolution. ∎
Remark 4.10.
Note that from the proof of Proposition 4.9 we can see that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality pair and $R$ is $n$-perfect, then $\mathcal{GF}_{\mathcal{A}}\subseteq{[\mathcal{GP}_{\mathcal{L}}]}^{\wedge}_{n}$.
When the ring $R$ is right coherent then the duality pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ es is symmetric, and since $(\mathcal{F},\mathcal{F}^{\bot})$ is always a complete cotorsion pair it follows that $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair. Thus, when $R$ is also left $n$-perfect we recover the main result in [13], as follows .
Proposition 4.11.
Let $R$ be a ring right coherent and left $n$-perfect. The following statements are true.
  • (a)
    The class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering in $\mathrm{Mod}(R)$
  • (b)
    The class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering in $\mathrm{Mod}(R)$.
Proof.
(a) Since $R$ is right coherent, the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair, then from Proposition 4.9 we get that $\mathcal{GP}_{\mathcal{F}}$ is special precovering, but this class is precisely the class of Ding projectives.
(b) Note that we have $\mathcal{P}(R)\subseteq\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$, thus from [4, Proposition 6.7] we get $\mathcal{GP}_{\mathcal{F}}=\mathcal{DP}(R)$, therefore the result it follows from (a). ∎
Since from [17, Proposition 23] the condition of being special precovering is sufficient to obtain a complete and hereditary cotorsion pair, we have the following result, which is a summary of what we have proved so far.
Proposition 4.12.
Let $R$ be a ring. The following statements are true.
  • (i)
    If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
  • (ii)
    If $R$ is left n-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$, then $(\mathcal{DP}(R),\mathcal{DP}^{\bot})=(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
  • (iii)
    If $(\mathcal{L},\mathcal{A})$ is a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect, then the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
Now we will see that it is possible to get more cotorsion pairs induced by these. In fact in each case of the above proposition we will obtain a family of cotorsion pairs. In order to obtain such cotorsion pairs we need the following lemma.
Lemma 4.13.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. For $M\in\mathrm{Mod}(R)$ and $m\in\mathbb{N}$ consider the following statements
  • (i)
    $M\in[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$,
  • (ii)
    $\mathrm{Ext}^{1}_{R}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$.
Then $\mathrm{(i)}\Rightarrow\mathrm{(ii)}$. If the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then $\mathrm{(ii)}\Rightarrow\mathrm{(i)}$, and thus both conditions are equivalent.
Proof.
We will use the following equality $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. From [5, Corollaries 5.2 (b) and 4.3 (c) ] we have the containment $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Now if $M\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ then from [5, Theorem 4.1 (b)] there is an exact sequence $\gamma:0\to M\to H\to G\to 0$ with $H\in\mathcal{P}(R)^{\wedge}_{m}$ and $G\in\mathcal{GP}_{\mathcal{L}}$, this implies that the previous sequence $\gamma$ splits, since $M\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, and so $M$ is direct summand of $H\in\mathcal{P}(R)^{\wedge}_{m}$.
$\mathrm{(i)}\Rightarrow\mathrm{(ii)}$. By induction over $m$. If $m=0$ is clear. Now assume that $m>0$. From [5, Theorem 4.1 (a)] there is an exact sequence $\theta:0\to K\to G\to M\to 0$ with $K\in\mathcal{P}(R)^{\wedge}_{m-1}$ and $G\in\mathcal{GP}_{\mathcal{L}}$. By definition there is an exact sequence $0\to G\to P\to G^{\prime}\to 0$ with $P\in\mathcal{P}(R)$ and $G^{\prime}\in\mathcal{GP}_{\mathcal{L}}$. We can construct the following p.o digram
[…]
Where $Q\in\mathcal{P}(R)^{\wedge}_{m}$. For $\overline{E}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$, applying $\mathrm{Hom}_{R}(-,\overline{E})$ we obtain the following commutative diagram
${\mathrm{Hom}_{R}(M,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(G,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(K,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(Q,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(P,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(K,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Ext}^{1}_{R}(Q,\overline{E})=0,}$
Thus $\mathrm{Hom}_{R}(G,\overline{E})\to\mathrm{Hom}_{R}(K,\overline{E})$ is an epimorphism for all $\overline{E}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Now assume that $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. From the above and $\theta$ we have the exact sequence
$\mathrm{Hom}_{R}(G,E)\twoheadrightarrow\mathrm{Hom}_{R}(K,E)\to\mathrm{Ext}^{1}_{R}(M,E)\to\mathrm{Ext}^{1}_{R}(G,E)=0$
Therefore, we conclude $\mathrm{Ext}_{R}^{1}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$.
$\mathrm{(ii)}\Rightarrow\mathrm{(i)}$. Let us suppose that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair and that for $M\in\mathrm{Mod}(R)$ we have $\mathrm{Ext}_{R}^{1}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. We will use the equivalence of [5, Corollary 4.3 (c)]. For $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to M\to H\to Q\to 0$, with $H\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ and $Q\in\mathcal{GP}_{\mathcal{L}}$. Applying $\mathrm{Hom}_{R}(-,E)$ we have the exact sequence
$\mathrm{Ext}_{R}^{1}(Q,E)\to\mathrm{Ext}_{R}^{1}(H,E)\to\mathrm{Ext}_{R}^{1}(M,E)=0$
where $\mathrm{Ext}_{R}^{1}(Q,E)=0$ since $Q\in\mathcal{GP}_{\mathcal{L}}$ and $E\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Therefore $\mathrm{Ext}_{R}^{1}(H,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, we will use this fact at the end.
From [10, Theorem 7.4.6], for $H$ there is an exact sequence $0\to K^{\prime}\to H^{\prime}\to H\to 0$, with $H^{\prime}\in\mathcal{P}(R)^{\wedge}_{m}$ and $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Note that $H,H^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ (since $\mathcal{P}(R)\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ implies $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ by the dual of [5, Lemma 2.6]). This implies that $K^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, i.e. $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, so that $\mathrm{Ext}^{1}_{R}(H,K^{\prime})=0$. That is $0\to K^{\prime}\to H^{\prime}\to H\to 0$ splits, therefore $H\in\mathcal{P}(R)^{\wedge}_{m}$. Thus, the exact sequence $0\to M\to H\to Q\to 0$ fulfils with the conditions of [5, Corollary 4.3 (c)]. ∎
With the above, we are ready to get a family of cotorsion pairs.
Theorem 4.14.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then for each $m>0$ the pair
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right),}$
is a complete and hereditary cotorsion pair.
Proof.
Form Lemma 4.13 we know that
$[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ =([$={{}^{\bot_{1}}\displaystyle{\left([\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right)}}$ and $[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ ([$\subseteq{{}^{\bot_{1}}([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m})}.$
We have for the other hand that $\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}=[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Then, taking orthogonals
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}\subseteq\left(\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}}.$
This implies that $\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Since the class $[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ is closed by kernels of epimorphisms [5, Corollary 4.10], we conclude that such cotorsion pair is hereditary. It remains to prove that such a pair is complete. We know that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is complete as well as $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ [10, Theorem 7.4.6], thus for $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to K\to H\to M\to 0\to 0$ with $H\in\mathcal{GP}_{\mathcal{L}}$ and $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. And to this $K$ there is $0\to K\to S\to E\to 0$ with $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}$. Consider the following p.o. diagram
[…]
Where $S\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, since $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. That is $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Now, since $H\in\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ it follows that $G\in[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Thus the exact sequence that works is $0\to S\to G\to M\to 0$. ∎
As a natural consequence we have a Hovey triple in $\mathrm{Mod}(R)$.
Corollary 4.15.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then for each $m>0$ there is a hereditary Hovey triple in $\mathrm{Mod}(R)$ given by
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;\mathcal{GP}_{\mathcal{L}}^{\bot_{1}},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\right)}$
Whose homotopy category is the stable category $[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}/([\mathcal{P}(R)^{\wedge}_{m}]\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$.
Proof.
We will apply [15, Theorem 1.2]. Note that the hereditary and complete cotorsion pairs $([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ and $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ are compatible, since from the proof of Lemma 4.13 we know that $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. While the last assertion follows from [22, Theorem 6.21]. ∎
Finally, to obtain a further application of the theory developed so far, we will address a duality pair that we believe to be very interesting.
Semidualizing bimodule
[2, 3] Consider $R$ and $S$ fixed associative rings with identities. An $(S,R)$-bimodule ${}_{S}C_{R}$ is called semidualizing if the following conditions are satisfied.
  • (a1)
    ${}_{S}C$ admits a degreewise finite $S$-projective resolution.
  • (a2)
    $C_{R}$ admits a degreewise finite $R^{\mathrm{op}}$-projective resolution.
  • (b1)
    The homothety map ${}_{S}S_{S}\to\mathrm{Hom}_{R^{\mathrm{op}}}(C,C)$ is an isomorphism.
  • (b2)
    The homothety map ${}_{R}R_{R}\to\mathrm{Hom}_{S}(C,C)$ is an isomorphism.
  • (c1)
    $\mathrm{Ext}_{S}^{\geq 1}(C,C)=0$.
  • (c2)
    $\mathrm{Ext}_{R^{\mathrm{op}}}^{\geq 1}(C,C)=0$.
Wakamatsu introduced in [24] and studied the named generalized tilting modules, usually called Wakamatsu tilting modules. Note that a bimodule ${}_{S}C_{R}$ is semidualizing if and only if it is Wakamatsu tilting [25, Corollary 3.2].
Associated to a semidualizing bimodule ${}_{S}C_{R}$ we have the Auslander and Bass classes.
(A) The Auslander class $\mathcal{A}_{C}(R)$ with respect to ${}_{S}C_{R}$ consists of all modules $M\in\mathrm{Mod}(R)$ satisfying the following conditions:
  • (A1)
    $\mathrm{Tor}^{R}_{\geq 1}(C,M)=0$,
  • (A2)
    $\mathrm{Ext}_{S}^{\geq 1}(C,C\otimes_{R}M)=0$,
  • (A3)
    The natural evaluation $\mu_{M}:M\to\mathrm{Hom}_{S}(C,C\otimes_{R}M)$ given by $\mu_{M}(x)(c)=c\otimes x$ for any $x\in M$ and $c\in C$, is an isomorphism in $\mathrm{Mod}(R)$.
(B) The Bass class $\mathcal{B}_{C}(S)$ with respect to ${}_{S}C_{R}$ consists of all modules $N\in\mathrm{Mod}(S)$ satisfying the following conditions:
  • (B1)
    $\mathrm{Ext}_{S}^{\geq 1}(C,N)=0$,
  • (B2)
    $\mathrm{Tor}_{\geq 1}^{R}(C,\mathrm{Hom}_{S}(C,N))=0$,
  • (B3)
    The natural evaluation $\nu_{N}:C\otimes_{R}\mathrm{Hom}_{S}(C,N)\to N$ given by $\nu_{N}(c\otimes f)=f(c)$ for any $c\in C$ and $f\in\mathrm{Hom}_{S}(C,N)$, is an isomorphism in $\mathrm{Mod}(S)$.
The Auslander class $\mathcal{A}_{C}(S^{\mathrm{op}})$ and the Bass class $\mathcal{B}_{C}(R^{\mathrm{op}})$ with respect to ${}_{S}C_{R}$ are defined similarly.
Recently been shown in [19] that there are two duality pairs [^5] Note that in the reference [19] the bimodule is ${}_{R}C_{S}$, and in our text is ${}_{S}C_{R}$ as appear in [3] associated to the bimodule ${}_{S}C_{R}$. Rewritten for the semidualizing bimodule ${}_{S}C_{R}$ (this is; in [19] we change $R$ by $S$) as follows.
By [19, Theorem 3.3] in $\mathrm{Mod}(S^{\mathrm{op}})$, the pair
$(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$
is a symmetric duality pair, and by [19, Theorem 3.3] in $\mathrm{Mod}(R)$ the pair
$(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$
is a symmetric duality pair. Furthermore, the pairs $(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$ and $(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$ are perfect duality pairs, respectively by [19, Corollaries 3.4 & 3.6]. From the above we have the following result, whose proof follows from Proposition 4.9.
Proposition 4.16.
Let ${}_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule, and consider the associated Auslader classes $\mathcal{A}_{C}(R)$ and $\mathcal{A}_{C}(S^{\mathrm{op}})$. The following statements are true.
  • (a)
    If $R$ is left n-perfect then the class of Gorenstein $\mathcal{A}_{C}(R)$-projectives, denoted by $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering over $\mathrm{Mod}(R)$.
  • (b)
    If $S^{\mathrm{op}}$ is left n-perfect then the class of Goresntein $\mathcal{A}_{C}(S^{\mathrm{op}})$-projectives $\mathcal{GP}_{\mathcal{A}_{C}}(S^{\mathrm{op}})$ is special precovering over $\mathrm{Mod}(S^{\mathrm{op}})$.
  • (c)
    If $R$ is left n-perfect and $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, then the class $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$
  • (d)
    If $S^{\mathrm{op}}$ is left n-perfect and $\mathcal{A}_{C}(S^{\mathrm{op}})\subseteq\mathcal{P}(S^{\mathrm{op}})^{\wedge}$, then the class $\mathcal{GP}(S^{\mathrm{op}})$ is special precovering on $\mathrm{Mod}(S^{\mathrm{op}})$.
Acknowledgements The author want to thank professor Raymundo Bautista (Centro de Ciencias Matemáticas - Universidad Nacional Autónoma de México) for several helpful discussions on the results of this article.
Funding The author was fully supported by a CONAHCyT posdoctoral fellowship at the Centro de Ciencias Matemáticas, UNAM.