### v3

# Some remarks on Gorenstein projective precovers

Abstract.
The existence of the Gorenstein projective precovers over arbitrary rings is an open question. In this paper, we make use of three different techniques addressing intrinsic and homological properties of several classes of relative Gorenstein projective $R$-modules, among them including the Gorenstein projectives and Ding projectives, with the purpose of giving some situations where Gorenstein projective precovers exists. Within the development of such techniques we obtaint a family of hereditary and complete cotorsion pairs and hereditary Hovey triples that comes from relative Gorenstein projective $R$-modules. We also study a class of Gorenstein projective $R$-modules relative to the Auslander class $\mathcal{A}_{C}(R)$ of a semidualizing $(R,S)$-bimodule $_{R}C_{S}$, where we make use of a property of “reduction”.
2010 Mathematics Subject Classification. Primary 18G10, 18G20, 18G25. Secondary 16E10.
Key Words: Gorenstein projective, Contravariantly finite, Hovey triple, Auslander class, Duality pair
1. Introduction
For $R$ an associative ring with identity, the Gorenstein projective, injective and flat $R$-modules where introduced in [14], since then the Gorenstein homological algebra has been developing intensively as a relative version of homological algebra that replaces the classical projective, injective and flat modules and resolutions with the Gorenstein versions. While in the classical homological algebra is know the existence of projective resolutions (resp. injective and flat resolutions) for any $R$-module without restrictions over $R$, the situation is different for the Gorenstein version. The question: What is the most general class of rings over which all modules have Gorenstein projective (injective) resolutions? still open. A comprehensive response has been given by S. Estrada, A. Iacob and K. Yeomans, who have proved that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$ provided that the ring $R$ be right coherent and left $n$-perfect [19, Theorem 2]. More recently, a relationship has been found between the existence of Gorenstein projective precovers and finitely presented $R$-modules with the Second Finitistic Dimension Conjeture [17, Theorem 5]. Furthermore, a closer relationship has been presented by P. Moradifar and J. Šaroch in [29] where is proved that contravariant finiteness of the class [^1] In this paper we use precovering as synonym of contravariantly finite. $\mathcal{GP}(R)^{<\infty}_{fin}$ (finitely generated $R$-modules of finite Gorenstein projective dimension) implies validity of the Second Finitistic Dimension Conjeture over left artinian rings. Furthermore is proved that that contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$ implies contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, over rings where $\mathcal{GP}(R)_{fin}$ (the class of finitely generated Gorenstein projective $R$-modules) is contravariantly finite and the converse holds for Artin algebras. Such relation is important, since was proved by Auslander and Reiten in [1] that contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, referred to as the Auslander–Reiten condition, is a sufficient condition for validity of the Second Finitistic Dimension Conjeture over an Artin algebra.
In view of the importance contravariant finiteness of the class $\mathcal{GP}(R)$, we study when such class is precovering on $\mathrm{Mod}(R)$. Although we know by M. Cortés-Izurdiaga, J. Šaroch [10] that the pair $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is always an hereditary cotorsion pair, a condition to be complete is that all projective modules are $\lambda$-pure-injective for some infinite regular cardinal $\lambda$ (in particular, if R is right $\Sigma$-pure-injective). In this paper we make use of three different ways to obtain $\mathcal{GP}(R)$ precovering. First, we make use of the intrinsic properties of $\mathcal{GP}(R)$ to see when the Holm’s question [13, Remark 4.5 (4)] is fulfilled. For the second, we make use of the different generalizations of the class $\mathcal{GP}(R)$ analyzing when these classes match. For the third, we make use of the technique of S. Estrada and A. Iacob [17], by finding a suitable complete hereditary cotorsion pair.
2. Preliminaries
In what follows, we shall work with categories of modules over an associative ring $R$ with identity. By $\mathrm{Mod}(R)$ and $\mathrm{Mod}(R^{\rm op})$ we denote the categories of left and right $R$-modules.
Projective, injective and flat $R$-modules will be important to present some definitions, remarks and examples. The classes of projective left and right $R$-modules will be denoted by $\mathcal{P}(R)$ and $\mathcal{P}(R^{\rm op})$, respectively. Similarly, we shall use the notations $\mathcal{I}(R)$, $\mathcal{I}(R^{\rm op})$, $\mathcal{F}(R)$ and $\mathcal{F}(R^{\rm op})$ for the classes of injective and flat modules in $\mathrm{Mod}(R)$ and $\mathrm{Mod}(R^{\rm op})$, respectively.
Concerning functors defined on modules, $\mathrm{Ext}^{i}_{R}(-,-)$ denotes the right $i$-th derived functor of $\mathrm{Hom}_{R}(-,-)$. If $M\in\mathrm{Mod}(R^{\rm op})$ and $N\in\mathrm{Mod}(R)$, $M\otimes_{R}N$ denotes the tensor product of $M$ and $N$. Recall the construction of this tensor products defines a bifunctor $\otimes_{R}-\colon\mathrm{Mod}(R^{\rm op})\times\mathrm{Mod}(R)\longrightarrow\mathrm{Mod}(\mathbb{Z})$, where $\mathrm{Mod}(\mathbb{Z})$ is the category of abelian groups.
Orthogonality
Let $\mathcal{X}\subseteq\mathrm{Mod}(R)$, $i\geq 1$ be a positive integer and $N\in\mathrm{Mod}(R)$. The expression $\mathrm{Ext}^{i}_{R}(\mathcal{X},N)=0$ means that $\mathrm{Ext}^{i}_{R}(X,N)=0$ for every $X\in\mathcal{X}$. Moreover, $\mathrm{Ext}^{i}_{R}(\mathcal{X,Y})=0$ if $\mathrm{Ext}^{i}_{R}(\mathcal{X},Y)=0$ for every $Y\in\mathcal{Y}$. The expression $\mathrm{Ext}^{i}_{R}(N,\mathcal{Y})=0$ has a similar meaning. Moreover, by $\mathrm{Ext}^{\geq 1}_{R}(M,N)=0$ we mean that $\mathrm{Ext}^{i}_{R}(M,N)=0$ for every $i\geq 1$. One also has similar meanings for $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X},N)=0$, $\mathrm{Ext}^{\geq 1}_{R}(N,\mathcal{Y})=0$ and $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$. We can also replace $\mathrm{Ext}$ by $\mathrm{Tor}$ in order to obtain the notations for $\mathrm{Tor}$-orthogonality. The right orthogonal complements of $\mathcal{X}$ will be denoted by
$\mathcal{X}^{\perp_{i}}$ $:=\{M\in\mathrm{Mod}(R){\rm\ :\ }\mathrm{Ext}^{i}_{R}(\mathcal{X},M)=0\}$ and $\mathcal{X}^{\perp}$ $:=\bigcap_{i\geq 1}\mathcal{X}^{\perp_{i}}.$
The left orthogonal complements, on the other hand, are defined similarly.
Relative homological dimensions
There are homological dimensions defined in terms of extension functors. Let $M\in\mathrm{Mod}(R)$ and $\mathcal{X},\mathcal{Y}\subseteq\mathrm{Mod}(R)$. The injective dimensions of $M$ and $\mathcal{Y}$ relative to $\mathcal{X}$ are defined by
$\mathrm{id}_{\mathcal{X}}(M)$ $:=\inf\{m\in\mathbb{Z}_{\geq 0}\text{ : }\mathrm{Ext}_{R}^{\geq m+1}(\mathcal{X},M)=0\}$ and $\mathrm{id}_{\mathcal{X}}(\mathcal{Y})$ $:=\sup\{\mathrm{id}_{\mathcal{X}}(Y)\text{ : }Y\in\mathcal{Y}\}.$
In the case where $\mathcal{X}=\mathrm{Mod}(R)$, we write
$\mathrm{id}_{\mathrm{Mod}(R)}(M)$ $=\mathrm{id}(M)$ and $\mathrm{id}_{\mathsf{Mod}(R)}(\mathcal{Y})$ $=\mathrm{id}(\mathcal{Y})$
for the (absolute) injective dimensions of $M$ and $\mathcal{Y}$. Dually we can define the relative dimensions $\mathrm{pd}_{\mathcal{X}}(M)$, $\mathrm{pd}_{\mathcal{X}}(\mathcal{Y})$ and $\mathrm{pd}(M)$, $\mathrm{pd}(\mathcal{Y})$.
By an $\mathcal{X}$-resolution of $M$ we mean an exact complex
$\cdots\to X_{m}\to X_{m-1}\to\cdots\to X_{1}\to X_{0}\to M\to 0$
with $X_{k}\in\mathcal{X}$ for every $k\in\mathbb{Z}_{\geq 0}$. If $X_{k}=0$ for $k>m$, we say that the previous resolution has length $m$. The resolution dimension relative to $\mathcal{X}$ (or the $\mathcal{X}$-resolution dimension) of $M$ is defined as the value
[…]
Moreover, if $\mathcal{Y}\subseteq\mathrm{Mod}(R)$ then
$\mathrm{resdim}_{\mathcal{X}}(\mathcal{Y}):=\sup\{\mathrm{resdim}_{\mathcal{X}}(Y)\ \mbox{ : }\ \text{$Y\in\mathcal{Y}$}\}$
defines the resolution dimension of $\mathcal{Y}$ relative to $\mathcal{X}$. The classes of objects with bounded (by some $n\geq 0$) and finite $\mathcal{X}$-resolution dimensions will be denoted by
$\mathcal{X}^{\wedge}_{n}$ $:=\{M\in\mathrm{Mod}(R)\text{ : }\mathrm{resdim}_{\mathcal{X}}(M)\leq n\}$ and $\mathcal{X}^{\wedge}$ $:=\bigcup_{n\geq 0}\mathcal{X}^{\wedge}_{n}.$
Dually, we can define $\mathcal{X}$-coresolutions and the coresolution dimension of $M$ and $\mathcal{Y}$ relative to $\mathcal{X}$ (denoted $\mathrm{coresdim}_{\mathcal{X}}(M)$ and $\mathrm{coresdim}_{\mathcal{X}}(\mathcal{Y})$). We also have the dual notations $\mathcal{X}^{\vee}_{n}$ and $\mathcal{X}^{\vee}$ for the classes of $R$-modules with bounded and finite $\mathcal{X}$-coresolution dimension.
Approximations
Given a class $\mathcal{X}$ of left $R$-modules and $M\in\mathrm{Mod}(R)$, recall that a morphism $\varphi\colon X\to M$ with $X\in\mathcal{X}$ is an $\mathcal{X}$-precover of $M$ if for every morphism $\varphi^{\prime}\colon X^{\prime}\to M$ with $X^{\prime}\in\mathcal{X}$, there exists a morphism $h\colon X^{\prime}\to X$ such that $\varphi^{\prime}=\varphi\circ h$. An $\mathcal{X}$-precover $\varphi$ is special if $\mathrm{CoKer}(\varphi)=0$ and $\mathrm{Ker}(\varphi)\in\mathcal{X}^{\bot_{1}}$.
A class $\mathcal{X}\subseteq\mathrm{Mod}(R)$ is (pre)covering if every left $R$-module has an $\mathcal{X}$-(pre)cover. Dually, one has the notions of (pre)envelopes and (pre)enveloping and special (pre)enveloping classes.
Cotorsion pairs
Two classes $\mathcal{X,Y}\subseteq\mathrm{Mod}(R)$ of left $R$-modules for a cotorsion pair $(\mathcal{X,Y})$ if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$.
A cotorsion pair $(\mathcal{X,Y})$ is:
  • Complete if $\mathcal{X}$ is special precovering, that is, for every $M\in\mathrm{Mod}(R)$ there is a short exact sequence $0\to Y\to X\to M\to 0$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$; or equivalently, if $\mathcal{Y}$ is special preenveloping.
  • Hereditary if $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$; or equivalently, if $\mathcal{X}$ is resolving (meaning that $\mathcal{X}$ is closed under extensions and kernels of epimorphisms, and contains the projective left $R$-modules) or $\mathcal{Y}$ is coresolving.
Note that if $(\mathcal{X,Y})$ is a hereditary cotorsion pair, then $\mathcal{X}={}^{\perp}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp}$.
Duality pairs
The notion of duality pair was introduced by Holm and Jørgensen in [24]. Two classes $\mathcal{L}\subseteq\mathrm{Mod}(R)$ and $\mathcal{A}\subseteq\mathrm{Mod}(R^{\rm op})$ of left and right $R$-modules form a duality pair $(\mathcal{L,A})$ in $\mathrm{Mod}(R)$ if:
  1. (1)
    $L\in\mathcal{L}$ if, and only if, $L^{+}:=\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Q}/\mathbb{Z})\in\mathcal{A}$.
  2. (2)
    $\mathcal{A}$ is closed under direct summands and finite direct sums.
One has a similar notion of duality pair in the case where $\mathcal{L}$ is a class of right $R$-modules, and $\mathcal{A}$ is a class of left $R$-modules.
A duality pair $(\mathcal{L,A})$ is called:
  • (co)product-closed if $\mathcal{L}$ is closed under (co)products.
  • perfect if it is coproduct closed, $\mathcal{L}$ is closed under extensions and contains $R$ (regarded as a left $R$-module).
  • complete if $(\mathcal{A},\mathcal{L})$ is also a duality pair and $(\mathcal{L},\mathcal{A})$ has all the properties required to be a perfect duality pair.
3. Gorenstein projective precovers
We recall that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ consist of cycles of exact complexes of left projective $R$-modules which remains exact after applying the functor $\mathrm{Hom}_{R}(-,P)$ for all $P\in\mathcal{P}(R)$. Also the class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$ consist of cycles of exact complexes of left flat $R$-modules which remains exact after applying the funtor $I\otimes_{R}$-, for all $I\in\mathcal{I}(R^{\mathrm{op}})$.
From Šaroch and Št’ovíchěk’s [31, Remark in p.24-25] we know now that the condition $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ is true if and only if $\mathcal{GP}(R)=\mathcal{P}\mathcal{GF}(R)$, which, in a way, answers Holm’s question [13, Remark 4.5 (4)]. Where the class here denoted $\mathcal{P}\mathcal{GF}(R)$ is the presented in [31, §4] called the class of projectively coresolved Gorenstein flat $R$-modules which consist of cycles of exact complexes of left projective $R$-modules that remains exact after applying the functor $(I\otimes_{R}-)$ for all $I\in\mathcal{I}(R^{\mathrm{op}})$. Now, from [31, Theorem 4.9] the class $\mathcal{P}\mathcal{GF}(R)$ is always a special precovering class. We declare this situation as follows.
Remark 3.1.
Let $R$ be a ring such that $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$, then the class $\mathcal{GP}(R)$ is special precovering.
From the previous result we can see that the conditions asked in [19, Theorem 1] there are more than needed. In what follows we address a variety of conditions that make $\mathcal{GP}(R)$ a class special precovering in $\mathrm{Mod}(R)$, for do this we use other kinds of relative projective Gorenstein modules and other particular notions that have recently appeared in the literature..
The statements (i)-(iv) in the following Proposition are basically equivalent to Holm’s question. Their proof it follows from [35, Theorem 2.4, Corollaries 2.5, 2,6] and Remark 3.1. It may be noted that there are other conditions involving the copure dimensions defined by Enochs and Jenda [14], which we do not deal with here.
Proposition 3.2.
Given a ring $R$ the class $\mathcal{GP}(R)$ is special precovering in each one of the following situations:
  • (i)
    $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$,
  • (ii)
    $\mathcal{GP}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}$,
  • (iii)
    $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)^{\wedge}$,
  • (iv)
    There exist an integer $n\geq 0$ such that $\mathrm{Tor}^{R}_{n}(I,M)=0$ for any $I\in\mathcal{I}(R^{\mathrm{op}})$ and $M\in\mathcal{GP}(R)$,
  • (v)
    If $R$ is either a right coherent ring such that $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ or a ring such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R)^{\wedge}$.
Note that a situation where $R$ is close to being right coherent and left $n$-perfect  [^2] This means that for the ring $R$ and $n\geq 0$ the containment $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$ is given. [19, Theorem 2] arises in Proposition 3.2 (v). In the following result we give a proof with another arguments where also $\mathcal{GP}(R)$ is special precovering. The purpose of giving the details is that the arguments can be adapted to other situations, where some kind of relative Gorenstein projective $R$-modules match with $\mathcal{GP}(R)$.
Proposition 3.3.
Let $R$ be a right coherent ring, then the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering on $\mathrm{Mod}(R)$. Furthermore, if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ then $\mathcal{GP}(R)$ is also special precovering on $\mathrm{Mod}(R)$.
Proof.
Indeed, since $R$ is right coherent, the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a symmetric duality pair, thus from [7, Theorem A.6.] [^3] Note that for D. Bravo et. al. the notion of duality pair differs from the one presented here. we get that $\mathcal{P}\mathcal{GF}(R)=\mathcal{DP}(R)$. Now from [31, Theorem 4.9] the class $\mathcal{P}\mathcal{GF}(R)$ is always a special precovering class, thus $\mathcal{DP}(R)$ is special precovering on $\mathrm{Mod}(R)$. Finally, if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$, we get from [4, Proposition 6.7] that $\mathcal{DP}(R)$ and $\mathcal{GP}(R)$ coincide, therefore $\mathcal{GP}(R)$ is special precovering. ∎
As is known, the class of Ding projectives appears as a generalization of Gorenstein projectives, while the class $\mathcal{P}\mathcal{GF}(R)$ comes to complement $\mathcal{GP}(R)$ since $\mathcal{P}\mathcal{GF}(R)$ is a subclass of $\mathcal{GP}(R)$ which consists of Gorenstein flat $R$-modules, which makes it more versatile, for example to show that every ring is GF-closed [31]. Thus, the generalizations and variants of $\mathcal{GP}(R)$ and $\mathcal{GF}(R)$ provide new information about them [28]. this is partly the reason why we address generalizations of $\mathcal{GP}(R)$ and $\mathcal{GF}(R)$ in order to provide a better understanding of them.
3.1. Precovers from generalized Gorenstein $R$-modules
In what follows, we shall consider classes $\mathcal{X}\subseteq\mathrm{Mod}(R)$ and $\mathcal{Y}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. The following definition of Gorenstein flat $R$-modules relative to a pair $(\mathcal{X},\mathcal{Y})$ comes from [36, Definition 2.1].
Definition 3.4.
An $R$-module $M$ is Gorenstein $(\mathcal{X},\mathcal{Y})$-flat if there exists an exact and $(\mathcal{Y}\otimes_{R}-)$-acyclic complex $X_{\bullet}\in\mathrm{Ch}(\mathcal{X})$ such that $M\cong Z_{0}(X_{\bullet})$. By $(\mathcal{Y}\otimes_{R}-)$-acyclic we mean that $Y\otimes_{R}X_{\bullet}$ is an exact complex of abelian groups for every $Y\in\mathcal{Y}$. The class of Gorenstein $(\mathcal{X,Y})$-flat $R$-modules will be denoted by $\mathcal{GF}_{(\mathcal{X,Y})}$.
There are several types of classes Gorenstein $(\mathcal{X},\mathcal{Y})$-flat that have been presented in recent literature. The class $\mathcal{P}\mathcal{GF}$ presentes above have been generalized and extensively studied by S. Estrada, A. Iacob and M. A. Pérez in [18, Definition 2.6]. This class is called projectively coresolved Gorenstein $\mathcal{B}$-flat and here denoted $\mathcal{P}\mathcal{GF}_{\mathcal{B}}$. In such paper also is studied the class $\mathcal{GF}_{\mathcal{B}}$ (in the notation of Definition 3.4 is $\mathcal{GF}_{(\mathcal{F},\mathcal{B})}$ where the subscript $\mathcal{F}$ denotes $\mathcal{F}(R)$) where $\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ is sometimes a semi-definable class [18, Definition 2.7].
In a near environment, is defined by J. Gillespie in [22] for a complete duality pair $(\mathcal{L},\mathcal{A})$ the class of Gorenstein $(\mathcal{L},\mathcal{A})$-flat $R$-modules, here denoted by $\mathcal{GF}_{\mathcal{A}}$, and is studied [22, §5] some of their dimensions and model structures. In such paper also is defined the class of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules here denoted by $\mathcal{GP}_{\mathcal{L}}$ and the class of Gorenstein $(\mathcal{L},\mathcal{A})$-injective $R$-modules here denoted by $\mathcal{GI}_{\mathcal{A}}$. We see that when $(\mathcal{L},\mathcal{A})$ is a duality pair the classes $\mathcal{GP}_{\mathcal{L}}$, $\mathcal{GI}_{\mathcal{A},}$, $\mathcal{GF}_{\mathcal{A}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ have an interesting interaction between them (see [4], [18], [22], [23], [34]). In fact ,the classes $\mathcal{GP}_{\mathcal{L}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ agree sometimes. This is true for a symmetric duality pair $(\mathcal{L},\mathcal{A})$, from [7, Theorem A.6.] since the class $\mathcal{L}$ is closed by pure quotients [24, Theorem 3.1]. This fact had been mentioned previously by J. Gillespie and A. Iacob in [23, Corollary 14]. We declare this result as follows.
Proposition 3.5.
Given a symmetric duality pair $(\mathcal{L},\mathcal{A})$ in $\mathrm{Mod}(R)$ the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ of projectively coresolved Gorenstein $\mathcal{A}$-flat $R$-modules coincide with the class $\mathcal{GP}_{\mathcal{L}}$ of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules.
When for a symmetric duality pair $(\mathcal{L},\mathcal{A})$ the class $\mathcal{A}$ is semi-definable then $\mathcal{GP}_{\mathcal{L}}$ is special precovering in the whole category $\mathrm{Mod}(R)$, since [18, Theorem 2.13] the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ is the left part of a complete and hereditary cotorsion pair. A duality pair with such property is the definable pair mentioned in [23, Example 12], denoted $(\langle\mathcal{F}(R)\rangle,\langle\mathcal{I}(R^{\mathrm{op}})\rangle)$. That is, the class of Gorenstein $(\mathcal{P}(R),\langle\mathcal{F}(R)\rangle)$-projective $R$-modules is special precovering. We state this as follows.
Proposition 3.6.
Consider the definable duality pair $(\langle\mathcal{F}(R)\rangle,\langle\mathcal{I}(R^{\mathrm{op}})\rangle)$ in $\mathrm{Mod}(R)$, then the class $\mathcal{GP}_{\langle\mathcal{F}(R)\rangle}$ is the left part of a complete and hereditary cotorsion pair.
For each $n\geq 2$ there is a duality pair $(\mathcal{FP}_{n}\mbox{-}Flat(R),\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}}))$, which consists of the $FP_{n}$-flat left R-modules and $FP_{n}$-injective right $R$-modules [8, Theorems 5.5 & 5.6]. The right part $\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}})$ is definable [18, Example 2.21. (3)]. The classes of relative Gorenstein $R$-modules that comes from these duality pairs have been studied in [28] by A. Iacob, proving in a manner similar to that used in Proposition 3.6 that $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ is the left part of a complete and hereditary cotorsion pair, and thus is a special precovering class. From here, we declare the following result.
Proposition 3.7.
If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given then the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering.
Proof.
Note that we have $\mathcal{P}(R)\subseteq\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [4, Proposition 6.7] the classes $\mathcal{GP}(R)$ and $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ are the same. Now from [28, Theorem 3.8] we get that $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair. ∎
Based on the above result we are interested in the study of the conditions over a duality pair $(\mathcal{L},\mathcal{A})$, that implies that $\mathcal{GP}_{\mathcal{L}}$ will be special precovering. This conditions will be different to the notion of semidefinable for the class $\mathcal{A}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. For do this, we study interactions between the classes of relative Gorenstein flat and relative Gorenstein projective associated to a duality pair $(\mathcal{L},\mathcal{A})$. We begin by recalling some facts and a definition adapted to our setting.
Definition 3.8 (GP-admissible pair).
[5, Definition 3.1]. A pair $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}(R)\times\mathrm{Mod}(R)$ is GP-admissible if satisfies the following conditions:
  1. (1)
    $\mathrm{Ext}^{\geq 1}_{R}(\mathcal{X,Y})=0$.
  2. (2)
    For every $A\in\mathrm{Mod}(R)$ there is an epimorphism $X\to A$ with $X\in\mathcal{X}$.
  3. (3)
    $\mathcal{X}$ and $\mathcal{Y}$ are closed under finite coproducts.
  4. (4)
    $\mathcal{X}$ is closed under extensions.
  5. (5)
    $\mathcal{X}\cap\mathcal{Y}$ is a relative cogenerator in $\mathcal{X}$, that is, for every $X\in\mathcal{X}$ there is an exact sequence $0\to X\to W\to X^{\prime}\to 0$ with $X^{\prime}\in\mathcal{X}$ and $W\in\mathcal{X}\cap\mathcal{Y}$.
Facts 3.9.
  • (i)
    We see from [18, Corollary 2.20] that for a class of right $R$-modules $\mathcal{B}$ such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{B}$ which satisfies that $\mathcal{GF}_{\mathcal{B}}$ is closed under extensions, then the pair $(\mathcal{GF}_{\mathcal{B}},\mathcal{GF}_{\mathcal{B}}^{\bot_{1}})$ is a hereditary and complete cotorsion pair.
  • (ii)
    Note that when $(\mathcal{L},\mathcal{A})$ is a perfect duality pair, then the containments $\mathcal{P}(R)\subseteq\mathcal{L}$, and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{A}$ are given. Furthermore the pair $(\mathcal{P}(R),\mathcal{L})$ is GP-admissible (see [4, §3]).
  • (iii)
    For a complete duality pair $(\mathcal{L},\mathcal{A})$ we know from [23, Corollary 43] that $(\mathcal{GF}_{\mathcal{A}},\mathcal{GF}_{\mathcal{A}}^{\bot_{1}})$ is a perfect cotorsion pair. In consequence the class $\mathcal{GF}_{\mathcal{A}}$ is closed by extensions and by $\mathrm{(i)}$ and $\mathrm{(ii)}$ also is an hereditary cotorsion pair.
We are ready to prove a useful result, which is a generalization of [19, Main Result].
Theorem 3.10.
Let $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}\times\mathrm{Mod}(R^{\mathrm{op}})$ be with the following properties.
  • (i)
    $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{Y}$.
  • (ii)
    The pair $(\mathcal{P}(R),\mathcal{X})$ is GP-admissible.
  • (iii)
    The class $\mathcal{GF}_{\mathcal{Y}}$ is closed by extensions.
Assume that the inclusions $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ are given. Then the class $\mathcal{GP}_{\mathcal{X}}$ is special precovering.
Proof.
Take $X\in\mathrm{Mod}(R)$. From Facts 3.9 (i), we have that $(\mathcal{GF}_{\mathcal{Y}},\mathcal{GF}_{\mathcal{Y}}^{\bot})$ is a hereditary and complete cotorsión pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}_{\mathcal{Y}}$ and $Y\in\mathcal{GF}_{\mathcal{Y}}^{\bot}$. From the containment $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}$, we get $\mathcal{GF}_{\mathcal{Y}}^{\bot}\subseteq\mathcal{GP}_{\mathcal{X}}^{\bot}$. Since $\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ and $(\mathcal{P}(R),\mathcal{X})$ is a GP-admissible pair, then from [5, Theorem 4.1 (a)] for $N\in\mathcal{GF}_{\mathcal{Y}}$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{GP}_{\mathcal{X}}$ and $W\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. We can construct the following p.b digram
[…]
and obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{GP}_{\mathcal{X}}^{\bot}$, this implies that $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. Therefore we obtain the exact sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{GP}_{\mathcal{X}}$ and $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. That is, $T\to X$ is a special $\mathcal{GP}_{\mathcal{X}}$-precover. ∎
As application we have the following.
Corollary 3.11.
Let $R$ be a ring such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$ and $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$. Then, the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering.
Proof.
We will verify the conditions of Theorem 3.10 with the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$. To this end note that $(\mathcal{P}(R),\mathcal{F}(R))$ is a GP-admissible pair and $\mathcal{GP}_{\mathcal{F}}$ is precisely the class of Ding-projective $R$-modules $\mathcal{DP}(R)$, while $\mathcal{GF}_{\mathcal{I}(R^{\mathrm{op}})}$ is the usual class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$. From [35, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ and always is true that $\mathcal{DP}(R)\subseteq\mathcal{GP}(R)$, therefore $\mathcal{DP}(R)\subseteq\mathcal{GF}(R)$. Also, from [11, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}$ and by [23, Corollary 14] the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{DP}(R)$ is always true. All this give us $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}\subseteq\mathcal{DP}(R)^{\wedge}$. ∎
Is know from [12, Theorem 3.8] and [20, Corollaries 4.5 and 4.6] that when $R$ is a Ding-Chen ring the class $\mathcal{DP}(R)$ of Ding projective $R$-modules is a class special precovering, the result above and Proposition 3.3 exhibits other conditions of when this occurs and shows the usefulness of Theorem 3.10.
Until now, from Proposition 3.5 we know that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality the containment $\mathcal{GP}_{\mathcal{L}}\subseteq\mathcal{GF}_{\mathcal{A}}$ is given. By Facts 3.9, if $(\mathcal{L},\mathcal{A})$ is perfect, then are fulfilled the conditions $\mathrm{(i),(ii)}$ in the previous Theorem 3.10. While if $(\mathcal{L},\mathcal{A})$ is complete then $\mathrm{(iii)}$ in Theorem 3.10 is also true. In the present generality, we declare the following result.
Theorem 3.12.
Let $(\mathcal{L},\mathcal{A})$ be a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect ring. Then, the class $\mathcal{GP}_{\mathcal{L}}$ is special precovering.
Proof.
Since $(\mathcal{L},\mathcal{A})$ is complete, also is symmetric and perfect. Thus, the conditions in Theorem 3.10 are fulfilled, except for the containment $\mathcal{GF}_{\mathcal{A}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{L}}$. To prove such a containment we must assume that $R$ is $n$-perfect, that is $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$.
Take $G\in\mathcal{GF}_{\mathcal{A}}$, then there is an exact complex $\mathbf{N}$ of flat $R$-modules $(\mathcal{A}\otimes_{R}-)$-acyclic with $G=Z_{0}(\mathbf{N})$. Consider a partial projective resolution of $\mathbf{N}$
$0\to\mathbf{C}\to\mathbf{P}_{n-1}\xrightarrow{d_{n-1}}\mathbf{P}_{n-2}\xrightarrow{d_{n-2}}\cdots\to\mathbf{P}_{1}\xrightarrow{d_{1}}\mathbf{P}_{0}\xrightarrow{d_{0}}\mathbf{N}\to 0,$
where $\mathbf{C}$ is not projective (but will be an exact complex). From this, for each $j$ we have the exact sequence $0\to C_{j}\to P_{n-1,j}\to\cdots\to P_{0,j}\to N_{j}\to 0$, with every $P_{i,j}\in\mathcal{P}(R)$, and since $N_{j}\in\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$ it follows that $C_{j}$ is projective for all $j$. Also we have the exact sequence $0\to\mathrm{Ker}(d_{0})\to\mathbf{P}_{0}\to\mathbf{N}\to 0$, with $\mathrm{Ker}(d_{0})$ exact, since $\mathbf{P}_{0}$ and $\mathbf{N}$ are exact. For each $A\in\mathcal{A}$ by assumption $A\otimes_{R}\mathbf{N}$ is acyclic, and since $\mathbf{P}_{0}$ is an projective complex then $A\otimes_{R}\mathbf{P}_{0}$ is acyclic, this implies that $A\otimes_{R}\mathrm{Ker}(d_{0})$ is acyclic. Repeating this procedure we obtain that $\mathbf{C}$ is an exact complex and $A\otimes_{R}\mathbf{C}$ is acyclic for all $A\in\mathcal{A}$. Therefore $\mathbf{C}$ is an exact complex of projectives $(\mathcal{A}\otimes_{R}-)$-acyclic, i.e. $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}$. Note that for each $j$ we have the exact sequence
$0\to Z_{j}(\mathbf{C})\to Z_{j}(\mathbf{P}_{n-1})\to\cdots Z_{j}(\mathbf{P}_{0})\to Z_{j}(\mathbf{N})\to 0$
with $Z_{j}(\mathbf{P}_{i})\in\mathcal{P}(R)$ for each $j$ and where $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}=\mathcal{GP}_{\mathcal{L}}$ by Proposition 3.5. Thus, for $j=0$ we get the desired $\mathcal{GP}_{\mathcal{L}}$-resolution. ∎
Remark 3.13.
Note that, from the proof of Theorem 3.12 we see that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality pair and $R$ is $n$-perfect, then $\mathcal{GF}_{\mathcal{A}}\subseteq{[\mathcal{GP}_{\mathcal{L}}]}^{\wedge}_{n}$.
When $R$ is a right coherent ring the duality pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ es is symmetric, and since $(\mathcal{F},\mathcal{F}^{\bot})$ is always a complete cotorsion pair, it follows that $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair. Thus, in the case when $R$ is also left $n$-perfect we recover from Theorem 3.12 the main result in [19].
Since the condition of being special precovering is enough to obtain a complete and hereditary cotorsion pair [23, Proposition 23], we give in the following result a summary in such terms of what we have proved so far.
Proposition 3.14.
Let $R$ be a ring. The following statements are true.
  • (i)
    If the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
  • (ii)
    For $R$ a right coherent ring, the pair $(\mathcal{DP}(R),\mathcal{DP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$. Furthermore if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is also an complete and hereditary cotorsion pair.
  • (iii)
    If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
  • (iv)
    If $R$ is left n-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$, then $(\mathcal{DP}(R),\mathcal{DP}^{\bot})=(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
  • (v)
    If $(\mathcal{L},\mathcal{A})$ is a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect, then the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ is a complete and hereditary cotorsion pair.
Proof.
(i) It follows from Remark 3.1, or equivalently by Proposition One-Main.
(ii) This is Proposition 3.3.
(iii) Comes from Proposition 3.7.
(iv) This is Corollary 3.11, where the equality is given from [5, Proposition 6.7].
(v) It follows from Theorem 3.12. ∎
An interesting phenomena comes from the existence of the cotorsion pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ since it is possible to get more cotorsion pairs induced by these. In fact, in each case of the proposition above, we will obtain a family of cotorsion pairs. In order to obtain such cotorsion pairs we need the following lemma.
Lemma 3.15.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. For $M\in\mathrm{Mod}(R)$ and $m\in\mathbb{N}$ consider the following statements
  • (i)
    $M\in[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$,
  • (ii)
    $\mathrm{Ext}^{1}_{R}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$.
Then $\mathrm{(i)}\Rightarrow\mathrm{(ii)}$. If the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then $\mathrm{(ii)}\Rightarrow\mathrm{(i)}$, and thus both conditions are equivalent.
Proof.
We will use the following equality $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. From [5, Corollaries 5.2 (b) and 4.3 (c) ] we have the containment $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Now if $M\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ then from [5, Theorem 4.1 (b)] there is an exact sequence $\gamma:0\to M\to H\to G\to 0$ with $H\in\mathcal{P}(R)^{\wedge}_{m}$ and $G\in\mathcal{GP}_{\mathcal{L}}$, this implies that the previous sequence $\gamma$ splits, since $M\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, and so $M$ is direct summand of $H\in\mathcal{P}(R)^{\wedge}_{m}$.
$\mathrm{(i)}\Rightarrow\mathrm{(ii)}$. By induction over $m$. If $m=0$ is clear. Now assume that $m>0$. From [5, Theorem 4.1 (a)] there is an exact sequence $\theta:0\to K\to G\to M\to 0$ with $K\in\mathcal{P}(R)^{\wedge}_{m-1}$ and $G\in\mathcal{GP}_{\mathcal{L}}$. By definition there is an exact sequence $0\to G\to P\to G^{\prime}\to 0$ with $P\in\mathcal{P}(R)$ and $G^{\prime}\in\mathcal{GP}_{\mathcal{L}}$. We can construct the following p.o digram
[…]
Where $Q\in\mathcal{P}(R)^{\wedge}_{m}$. For $\overline{E}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$, applying $\mathrm{Hom}_{R}(-,\overline{E})$ we obtain the following commutative diagram
${\mathrm{Hom}_{R}(M,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(G,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(K,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(Q,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(P,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{R}(K,\overline{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Ext}^{1}_{R}(Q,\overline{E})=0,}$
Thus $\mathrm{Hom}_{R}(G,\overline{E})\to\mathrm{Hom}_{R}(K,\overline{E})$ is an epimorphism for all $\overline{E}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Now assume that $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. From the above and $\theta$ we have the exact sequence
$\mathrm{Hom}_{R}(G,E)\twoheadrightarrow\mathrm{Hom}_{R}(K,E)\to\mathrm{Ext}^{1}_{R}(M,E)\to\mathrm{Ext}^{1}_{R}(G,E)=0$
Therefore, we conclude $\mathrm{Ext}_{R}^{1}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$.
$\mathrm{(ii)}\Rightarrow\mathrm{(i)}$. Let us suppose that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair and that for $M\in\mathrm{Mod}(R)$ we have $\mathrm{Ext}_{R}^{1}(M,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. We will use the equivalence of [5, Corollary 4.3 (c)]. For $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to M\to H\to Q\to 0$, with $H\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ and $Q\in\mathcal{GP}_{\mathcal{L}}$. Applying $\mathrm{Hom}_{R}(-,E)$ we have the exact sequence
$\mathrm{Ext}_{R}^{1}(Q,E)\to\mathrm{Ext}_{R}^{1}(H,E)\to\mathrm{Ext}_{R}^{1}(M,E)=0$
where $\mathrm{Ext}_{R}^{1}(Q,E)=0$ since $Q\in\mathcal{GP}_{\mathcal{L}}$ and $E\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Therefore $\mathrm{Ext}_{R}^{1}(H,E)=0$ for all $E\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, we will use this fact at the end.
From [15, Theorem 7.4.6], for $H$ there is an exact sequence $0\to K^{\prime}\to H^{\prime}\to H\to 0$, with $H^{\prime}\in\mathcal{P}(R)^{\wedge}_{m}$ and $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Note that $H,H^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ (since $\mathcal{P}(R)\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ implies $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ by the dual of [5, Lemma 2.6]). This implies that $K^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, i.e. $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, so that $\mathrm{Ext}^{1}_{R}(H,K^{\prime})=0$. That is $0\to K^{\prime}\to H^{\prime}\to H\to 0$ splits, therefore $H\in\mathcal{P}(R)^{\wedge}_{m}$. Thus, the exact sequence $0\to M\to H\to Q\to 0$ fulfils with the conditions of [5, Corollary 4.3 (c)]. ∎
With the above, we are ready to get a family of cotorsion pairs.
Theorem 3.16.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then for each $m>0$ the pair
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right),}$
is a complete and hereditary cotorsion pair.
Proof.
From Lemma 3.15 we know that following;
$[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ $={{}^{\bot_{1}}\displaystyle{\left([\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right)}}$ and $[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ $\subseteq{([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m})^{\bot_{1}}}.$
For the other hand we have that $\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}=[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Then, taking orthogonals
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}\subseteq\left(\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}}.$
This implies that $\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Since the class $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ is closed by kernels of epimorphisms [5, Corollary 4.10], we conclude that such cotorsion pair is hereditary. It remains to prove that such a pair is complete. We know that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is complete as well as $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ [15, Theorem 7.4.6], thus for $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to K\to H\to M\to 0\to 0$ with $H\in\mathcal{GP}_{\mathcal{L}}$ and $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. And to this $K$ there is $0\to K\to S\to E\to 0$ with $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}$. Consider the following p.o. diagram
[…]
Where $S\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, since $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. That is $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Now, since $H\in\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ it follows that $G\in[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Thus the exact sequence that works is $0\to S\to G\to M\to 0$. ∎
Thus, in each of the conditions of Proposition 3.14, we get a family of cotorsion pairs. As an expected consequence we also obtain a Hovey triple in $\mathrm{Mod}(R)$.
Corollary 3.17.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair, then for each $m>0$ there is a hereditary Hovey triple in $\mathrm{Mod}(R)$ given by
${\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;\mathcal{GP}_{\mathcal{L}}^{\bot_{1}},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\right)}$
Whose homotopy category is the stable category $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}/([\mathcal{P}(R)^{\wedge}_{m}]\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$.
Proof.
We will apply [21, Theorem 1.2]. Note that the hereditary and complete cotorsion pairs $([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ and $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ are compatible, since from the proof of Lemma 3.15 we know that $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. While the last assertion follows from [30, Theorem 6.21]. ∎
As stated in the introduction, we are interested in the contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$,  [^4] Note that $\mathcal{GP}(R)_{\infty}$ in the notation of [29] refers to $\mathcal{GP}(R)^{\wedge}=\mathcal{GP}(R)^{<\infty}$ in this paper. since this implies the contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$ (such property implies the second finitistic dimension conjeture over an Artin algebra) over rings where $\mathcal{GP}(R)_{fin}$ is contravariantly finite. We will prove a kind of converse over $\mathrm{Mod}(R)$, in the generality of assuming that $\mathcal{GP}_{\mathcal{L}}$ is special precovering. We specify these ideas in the following result. But first we need to state a few notions adapted to our setting.
Definition 3.18.
[5, Definition 4.16] The finitistic $(\mathcal{P}(R),\mathcal{L})$-Gorenstein proyective dimension of $\mathrm{Mod}(R)$ is defined and denoted by
$\mathrm{FGPD}_{(\mathcal{P},\mathcal{L})}(R):=\sup\{\mathrm{Gpd}_{\mathcal{L}}(M):M\in\mathcal{GP}_{\mathcal{L}}^{\wedge}\}.$
where $\mathrm{Gpd}_{\mathcal{L}}(M):=\mathrm{resdim}_{\mathcal{GP}_{\mathcal{L}}}(M)$.
The finitistic projective dimension of $\mathrm{Mod}(R)$ is $\mathrm{FPD}(R):=\sup\{\mathrm{pd}(M):M\in\mathcal{P}(R)^{\wedge}\}$.
Note that the following result have a version over each case of Proposition 3.14.
Proposition 3.19.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair and $\mathrm{FPD}(R)=t<\infty$ then $\mathcal{GP}_{\mathcal{L}}^{<\infty}$ is the left hand side of a complete and hereditary cotorsion pair.
Proof.
Indeed, since $\mathrm{FPD}(R)=t$, from [5, Theorem 4.23 (a)] we get that $\mathrm{FGPD}_{(\mathcal{P},\mathcal{L})}(R)=t<\infty$. This implies from Theorem 3.16 that $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{t}=\mathcal{GP}_{\mathcal{L}}^{<\infty}$ is the left hand side of a complete and hereditary cotorsion pair. ∎
In the following, we compare the results obtained in this work with the second main result of P. Moradifar and J. Šaroch [29, (2.5) Theorem].
Remark 3.20.
Let us consider $\Lambda$ an Artin algebra. We know from [6, X, Theorem 2.4 (iv)] that $\mathcal{GP}(\Lambda)$ is the left hand side of a complete and hereditary cotorsion pair in $\mathrm{Mod}(\Lambda)$, so $\mathcal{GP}(\Lambda)$ is contravariantly finite in $\mathrm{Mod}(\Lambda)$. The following statements are true.
  • (i)
    For all $n\geq 0$, the class $\mathcal{GP}(\Lambda)^{\wedge}_{n}$ es contravariantly finite (compare with [29, (2.5) Theorem (i)]).
  • (ii)
    If the class $\mathcal{P}(\Lambda)^{<\infty}_{fin}$ of $\Lambda$-modules finitely generated of finite projective dimensiones contravariantly finite, by Huisgen-Zimmermann and Smalø [27], we get that $\mathrm{FPD}(\Lambda)<\infty$. Thus, from Proposition 3.19 we conclude that $\mathcal{GP}(\Lambda)^{\wedge}$ is contravariantly finite (compare with [29, (2.5) Theorem (ii)]).
Finally, to obtain a further application of the theory developed so far, we will address a duality pair from which we will be able to define a interesting class of Gorenstein $R$-modules on which we will apply our theory.
Semidualizing bimodule
[2, 25] Consider $R$ and $S$ fixed associative rings with identities. An $(S,R)$-bimodule $_{S}C_{R}$ is called semidualizing if the following conditions are satisfied.
  • (a1)
    $_{S}C$ admits a degreewise finite $S$-projective resolution.
  • (a2)
    $C_{R}$ admits a degreewise finite $R^{\mathrm{op}}$-projective resolution.
  • (b1)
    The homothety map $_{S}S_{S}\to\mathrm{Hom}_{R^{\mathrm{op}}}(C,C)$ is an isomorphism.
  • (b2)
    The homothety map $_{R}R_{R}\to\mathrm{Hom}_{S}(C,C)$ is an isomorphism.
  • (c1)
    $\mathrm{Ext}_{S}^{\geq 1}(C,C)=0$.
  • (c2)
    $\mathrm{Ext}_{R^{\mathrm{op}}}^{\geq 1}(C,C)=0$.
Wakamatsu introduced in [32] and studied the named generalized tilting modules, usually called Wakamatsu tilting modules. Note that a bimodule $_{S}C_{R}$ is semidualizing if and only if it is Wakamatsu tilting [33, Corollary 3.2].
Associated to a semidualizing bimodule $_{S}C_{R}$ we have the Auslander and Bass classes.
(A) The Auslander class $\mathcal{A}_{C}(R)$ with respect to $_{S}C_{R}$ consists of all modules $M\in\mathrm{Mod}(R)$ satisfying the following conditions:
  • (A1)
    $\mathrm{Tor}^{R}_{\geq 1}(C,M)=0$,
  • (A2)
    $\mathrm{Ext}_{S}^{\geq 1}(C,C\otimes_{R}M)=0$,
  • (A3)
    The natural evaluation $\mu_{M}:M\to\mathrm{Hom}_{S}(C,C\otimes_{R}M)$ given by $\mu_{M}(x)(c)=c\otimes x$ for any $x\in M$ and $c\in C$, is an isomorphism in $\mathrm{Mod}(R)$.
(B) The Bass class $\mathcal{B}_{C}(S)$ with respect to $_{S}C_{R}$ consists of all modules $N\in\mathrm{Mod}(S)$ satisfying the following conditions:
  • (B1)
    $\mathrm{Ext}_{S}^{\geq 1}(C,N)=0$,
  • (B2)
    $\mathrm{Tor}_{\geq 1}^{R}(C,\mathrm{Hom}_{S}(C,N))=0$,
  • (B3)
    The natural evaluation $\nu_{N}:C\otimes_{R}\mathrm{Hom}_{S}(C,N)\to N$ given by $\nu_{N}(c\otimes f)=f(c)$ for any $c\in C$ and $f\in\mathrm{Hom}_{S}(C,N)$, is an isomorphism in $\mathrm{Mod}(S)$.
The Auslander class $\mathcal{A}_{C}(S^{\mathrm{op}})$ and the Bass class $\mathcal{B}_{C}(R^{\mathrm{op}})$ with respect to $_{S}C_{R}$ are defined similarly.
Recently been shown in [26] that there are two duality pairs [^5] Note that in [26] the bimodule is $_{R}C_{S}$, and in our text is $_{S}C_{R}$ as appear in [25] associated to the bimodule $_{S}C_{R}$. Rewritten for the semidualizing bimodule $_{S}C_{R}$ (this is; in [26] we change $R$ by $S$) as follows.
By [26, Theorem 3.3] in $\mathrm{Mod}(S^{\mathrm{op}})$, the pair
$(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$
is a symmetric duality pair, and by [26, Theorem 3.3] in $\mathrm{Mod}(R)$ the pair
$(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$
is a symmetric duality pair. Furthermore, the pairs $(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$ and $(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$ are perfect duality pairs, respectively by [26, Corollaries 3.4 & 3.6]. From the above, we have the following result, whose proof follows from Proposition 3.12.
Proposition 3.21.
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule, and consider the associated Auslander class $\mathcal{A}_{C}(R)$. If $R$ is left n-perfect then the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ of Gorenstein $\mathcal{A}_{C}(R)$-projectives is special precovering. If in addition $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, then the class $\mathcal{GP}(R)$ is also special precovering.
Finally in (ii) of the following result, we will make use of a different technique that can also be used on other versions of Gorenstein projective $R$-modules, we refer to a kind of “reduction”, as in [17]. We will see that as an advantage of this technique, we can dispense the condition of being n-perfect in the above Proposition. For do this we recall the notion of $C$-injective $R$-modules [25, Definition 5.1] , which is the class $\mathcal{I}_{C}(R)$ consisting by $R$-modules of the form $\mathrm{Hom}_{S}(C,I)$ where $I\in\mathcal{I}(S)$.
Theorem 3.22.
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule. The class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering under the following situations.
  • (i)
    When $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$.
  • (ii)
    If $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$.
Proof.
(a) From [16, Proposition 3.9], we actually know that $\mathcal{GP}(R)\subseteq\mathcal{A}_{C}(R)$, in consequence $\mathcal{GP}_{\mathcal{A}_{C}}(R)\subseteq\mathcal{A}_{C}(R)$ since $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)$ from [25, Theorem 6.2]. This give us the containment $\mathcal{GP}_{\mathcal{A}_{C}}(R)\subseteq\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$, which implies that all $N\in\mathcal{A}_{C}(R)$ possesses a special $\mathcal{GP}_{\mathcal{A}_{C}}(R)$-precover, for the above mentioned, Facts 3.9 and [5, Theorem 4.1 (a)]. Also was recently proved that for a general ring the pair $(\mathcal{A}_{C}(R),\mathcal{A}_{C}(R)^{\bot})$ is a perfect cotorsion pair, thus, by [17, Theorem 1] we get that the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering.
(b) Now let us suppose that $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$, we will prove under this conditions that $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}_{m}$, and so the result will follow from (a). To do this, take $M\in\mathcal{A}_{C}(R)$. By [16, Proposition 3.12] there exists an exact sequence
$0\to M\to U^{0}\to U^{1}\to\cdots$
where each $U^{i}\in\mathcal{I}_{C}(R)$.
From the inequality $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and [9, Ch. XVII, §1 Proposition 1.3] we can construct the following commutative diagram
[…]
such that $P_{i}^{j}\in\mathcal{P}(R)$ for all $i\in\{0,1,\dots,m\}$ and all $j\geq 0$. With $Q_{i}:=\mathrm{Ker}(P^{0}_{i}\to P^{1}_{i})\in\mathcal{P}(R)$ for all $i\in\{0,\dots,m-1\}$ and
$\Lambda:0\to Q\to P_{m}^{0}\to P_{m}^{1}\to\cdots$
$\Theta:0\to Q\to Q_{m-1}\to\cdots\to Q_{0}\to M\to 0$
exact complexes. Note that the complex $\Lambda$ can be completed on the left with a resolution by projective $R$-modules, and since $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$, by [3, Lemma 3.6] such exact complex is $\mathrm{Hom}_{R}(-,\mathcal{A}_{C}(R))$-exact, from where $Q\in\mathcal{GP}_{\mathcal{A}_{C}}(R)$. But then, from the exact complex $\Theta$ we obtain that $\mathrm{resdim}_{\mathcal{GP}_{\mathcal{A}_{C}}(R)}(M)\leq m$, that is $M\in\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}_{m}$. ∎
The previous result allows us to obtain another family of cotorsion pairs and a family Hovey triples similarly to the Corollary 3.17. By another hand, the following result address the conditions of when the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ coincides with the classical $\mathcal{GP}(R)$.
Corollary 3.23.
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule and assume that $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$. Then, the class $\mathcal{GP}(R)$ is special precovering.
Proof.
Indeed, the condition $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$ implies that $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$, since by [25, Theorem 6.2] $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)$. Thus, from Theorem 3.22 (a), the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering. Currently, we also have the containments $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [5, Proposition 6.7] we obtain the equality $\mathcal{GP}_{\mathcal{A}_{C}}(R)=\mathcal{GP}(R)$. ∎
Acknowledgements The author want to thank professor Raymundo Bautista (Centro de Ciencias Matemáticas - Universidad Nacional Autónoma de México) for several helpful discussions on the results of this article.
Funding The author was fully supported by a CONAHCyT (actually renamed SECIHTI) posdoctoral fellowship CVU 443002, managed by the Universidad Michoacana de San Nicolas de Hidalgo at the Centro de Ciencias Matemáticas, UNAM.
Declarations
Ethical Approval Not applicable.
Availability of data and materials Not applicable.
Competing interests The authors declare no competing interests.